OpenStudy (twolsey5):

the equation 0=12+(-12) exemplifies which property

OpenStudy (twolsey5):

associative property of adition commutative property of addition miltiplicative inverse property additive inverse property

OpenStudy (abb0t):

additive inverse. because you're adding, but actually subtracting. Because you have 12 plus a negative number. You have 12 dollars, but you owe me 12. So you have 0 dollars left.

OpenStudy (zzr0ck3r):

associative says a+(b+c) = (a+b) +c commutative says a+b=b+a multiplicative inverse says a*a^{-1}=1 additive inverse says a+(-a)=0

OpenStudy (twolsey5):

thank you

OpenStudy (twolsey5):

1. The expression 12x plus 15x minus 11xcan be combined because of which property? /(1 point)/ * Commutative Property of Addition * Distributive Property * Additive Inverse Property * Multiplicative Inverse Property